OpenStudy (anonymous):

A right cylinder has a height of 5 inches and a radius of 7 inches. What is its volume in terms of ?? A. 35 ? cubic inches B. 175 ? cubic inches C. 245 ? cubic inches D. 12 ? cubic inches E. 875 ? cubic inches

Directrix (directrix):

C. 245 pi cubic inches V = Bh where B is the area of the circular base and h is the height V = pi r^2 h V = pi (7)^2 (5) V = 245 pi cubic inches